• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Yuan,Jing

    Log entry time 23:52:57 on August18,2003

    Entry number 111865

    keyword=swing shift summary

    We have no beam through out the whole swing shift ;

    We did septum train ,427K quench.after my calculation, the septum can work after 1:00am (Aug/19) if rate of temperature dropping is stable.

    Vince change the rate of septum current inceasing to 0.012V on 19:38

    MCC called CHL still tripped on 23:52 pm ,They cann't give the beam back on 9:00pm as they said before .They don't know when beam will be back.